lindsaylynn1120:

i need helppp A system of equations is given. −5y = 10 − 5x −2y = 8 − 4x Solve for (x, y) using the elimination method. Show all work

Vocaloid:

using elimination, you multiply or divide the equations by constants until you have terms you can eliminate through adding or subtracting the equations. this is a little easier to explain with the actual values in the equation. −5y = 10 − 5x −2y = 8 − 4x in general, you can multiply both sides of an equation by a constant and keep both sides equal. so in this case, you could multiply the first equation by 2 and the second equation by 5. −10y = 20 − 10x −10y = 40 − 20x notice how I've chosen the values to make both y terms -10y. that way, you can subtract the two equations and *eliminate* the y term, hence the method being called elimination. from there, you can solve for x. finally, pick any of the original equations, plug in the x-value, and solve for y.